OpenStudy (anonymous):

Which is the equation of the line passing through (0, 6) and (3, 0)?

OpenStudy (anonymous):

@iPwnBunnies @ikatouni @mlb004 @REALEST @random231

OpenStudy (anonymous):

@JayBaybie @rodrigorodriguez13

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

Are there choices?

OpenStudy (anonymous):

oh.. yeah, hang on!! lolz

OpenStudy (anonymous):

ok sorry im here

OpenStudy (anonymous):

A.) y= -1/2x + 6 B.) y= -1/2x + 3 C.) y= 2x +6 D.) y= 2x +3

OpenStudy (anonymous):

m = (y2 - y1)/(x2 - x1) m = (0 - 4)/(6 + 2) m = -4/8 m = -1/2 y - y1 = m(x - x1) y - 0 = -1/2(x - 6) y = (-1/2)x + 3 Answer B

OpenStudy (anonymous):

is he right?? @undeadknight26

undeadknight26 (undeadknight26):

I believe so...

OpenStudy (anonymous):

thanks!!!